A body of mass 1 kg thrown upwards with a velocity of 10 m/s comes to rest (momentarily) after moving 4m?


what is the work done by the air drag in the process

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Raghava,

m=mass=1kg, mv2/2=mgh+ work done by air drag

work done by air drag=1*(50-4g)=50-40=10Joule

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Raghava !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

Nanda RJ
12 years ago

potential energy=mgh==1*10*4=40

kinetic energy=1/2*m*v2  =50

work done by the air drag= sum of kinetic and potential energy=50+40=90Nm.